OpenStudy (anonymous):

4x+3y=9 8x+3y=-15 elimination method answer in ordered pairs

OpenStudy (saifoo.khan):

x = -6 y = 11

OpenStudy (anonymous):

4x+3y=9 is equation no 1

OpenStudy (anonymous):

8x+3y=-15 is equation 2 if we subtract 2 from 1 so we get 4x+3y=9 -(8x+3y=-15) -4x+0= 24 x=-6 4(-6)+3y=9 3y=9+24 y=33/3=11=y